# Unleashing Reasoning Capability of LLMs via Scalable Question Synthesis from Scratch

### v1

@@ Appendix A Additional Data Statistics
Appendix B Prompts
Prompts for Problem Solvability Optimization Please act as a professional math teacher. Your goal is to create high quality math word problems to help students learn math. You will be given a math question. Please optimize the Given Question and follow the instructions. To achieve the goal, please follow the steps: # Please check that the given question is a math question and write detailed solution to the Given Question. # Based on the problem-solving process, double check the question is solvable. # If you feel that the given question is not a meaningful math question, rewrite one that makes sense to you. Otherwise, modify the Given question according to your checking comment to ensure it is solvable and of high quality. # If the question can be solved with just a few simple thinking processes, you can rewrite it to explicitly request multiple-step reasoning. You have five principles to do this: # Ensure the optimized question only asks for one thing, be reasonable and solvable, be based on the Given Question (if possible), and can be answered with only a number (float or integer). For example, DO NOT ask, ‘what is the amount of A, B and C?’. # Ensure the optimized question is in line with common sense of life. For example, the amount someone has or pays must be a positive number, and the number of people must be an integer. # Ensure your student can answer the optimized question without the given question. If you want to use some numbers, conditions or background in the given question, please restate them to ensure no information is omitted in your optimized question. # Please DO NOT include solution in your question. Given Question: problem Your output should be in the following format: CREATED QUESTION: [your created question] VERIFICATION AND MODIFICATION: [solve the question step-by-step and modify it to follow all principles] FINAL QUESTION: [your final created question]
Figure 9: The prompts used to optimize the solvability of questions for QPO Training.
@@ Appendix B Prompts
Prompts for Problem Difficulty Optimization You are an Math Problem Rewriter that rewrites the given #Problem# into a more complex version. Please follow the steps below to rewrite the given ”#Problem#” into a more complex version. Step 1: Please read the ”#Problem#” carefully and list all the possible methods to make this problem more complex (to make it a bit harder for well-known AI assistants such as ChatGPT and GPT4 to handle). Note that the problem itself might be erroneous, and you need to first correct the errors within it. Step 2: Please create a comprehensive plan based on the #Methods List# generated in Step 1 to make the #Problem# more complex. The plan should include several methods from the #Methods List#. Step 3: Please execute the plan step by step and provide the #Rewritten Problem#. #Rewritten Problem# can only add 10 to 20 words into the ”#Problem#”. Step 4: Please carefully review the #Rewritten Problem# and identify any unreasonable parts. Ensure that the #Rewritten Problem# is only a more complex version of the #Problem#. Just provide the #Finally Rewritten Problem# without any explanation and step-by-step reasoning guidance. Please reply strictly in the following format: Step 1 #Methods List#: Step 2 #Plan#: Step 3 #Rewritten Problem#: Step 4 #Finally Rewritten Problem#: #Problem#: Problem
Figure 10: The prompts used to optimize the difficulty of questions for QPO Training.
@@
Prompts for Problem Solvability Check Please act as a professional math teacher. Your goal is to determine if the given problem is a valuable math problem. You need to consider two aspects: 1. The given problem is a math problem. 2. The given math problem can be solved based on the conditions provided in the problem (You can first try to solve it and then judge its solvability). Please reason step by step and conclude with either ‘Yes’ or ‘No’. Given Problem: Problem
Figure 11: The prompts used to check the solvability of questions.
Prompts for Difficulty Classification # Instruction You first need to identify the given user intent and then label the difficulty level of the user query based on the content of the user query. ## User Query ‘‘‘ Input ‘‘‘ ## Output Format Given the user query, in your output, you first need to identify the user intent and the knowledge needed to solve the task in the user query. Then, rate the difficulty level of the user query as very easy, easy, medium, hard, or very hard. Now, please output the user intent and difficulty level below in a json format by filling in the placeholders in []: ‘‘‘ {{ “intent”: “The user wants to [….]”, “knowledge”: “To solve this problem, the models need to know [….]”, “difficulty”: “[very easy/easy/medium/hard/very hard]” }} ‘‘‘
@@
Prompts for Topic Classification As a mathematics education specialist, please analyze the topics of the provided question and its answer. Specific requirements are as follows: 1. You should identify and categorize the main mathematical topics involved in the problem. If knowledge from non-mathematical fields is used, it is classified into Others - xxx, such as Others - Problem Context. 2. You should put your final answer between $<$TOPIC$>$ and $<$/TOPIC$>$. —— Question: Compute $\cos 330^{\circ}$. Answer: We know that $330^{\circ}=360^{\circ}-30^{\circ}$. Since $\cos(360^{\circ}-\theta)=\cos\theta$ for all angles $\theta$, we have $\cos 330^{\circ}=\cos 30^{\circ}$. Since $\cos 30^{\circ}=\frac{\sqrt{3}}{2}$, we can conclude that $\cos 330^{\circ}=\boxed{\frac{\sqrt{3}}{2}}$. Analysis: $<$TOPIC$>$Trigonometry - Cosine Function$<$/TOPIC$>$ —— Question: Question Answer: Answer Analysis:
@@
Examples for Solvability Optimization Problems 1 (Before Optimization): There are 10 survivors in an emergency room. Each survivor is either a child, a woman, or a man. If there are 4 men and 3 times as many women as men, how many children are there? Problems 1 (After Optimization): There are 10 survivors in an emergency room. Each survivor is either a child, a woman, or a man. If there are 4 men and an equal number of women as men, how many children are there?   Problems 2 (Before Optimization): How many sides does a polygon have if it is a regular polygon? Problems 2 (After Optimization): How many sides does a regular polygon have if each interior angle is 120 degrees?   Problems 3 (Before Optimization): Find the sum of the first three terms of this series. Problems 3 (After Optimization): Calculate the sum of the first three terms of the arithmetic series where the first term is 5 and the common difference is 3.
@@
Examples for Difficulty Optimization Problems 1 (Before Optimization): How many 4-digit positive integers are there? Problems 1 (After Optimization): How many 4-digit positive integers can be formed using non-repeating digits where the sum of these digits must be even, and the integers fall within the range of 1000 to $9$?   Problems 2 (Before Optimization): The average of 15 numbers is 32. An additional number is then added to the list, and the new average of the 16 numbers is 34. What number was added to the list? Problems 2 (After Optimization): The average of 15 positive integers is 32, but one integer fluctuates to 30 before adding a new number. After adding this new number, the average of the 16 integers becomes 34. Calculate the added number and find the standard deviation of all 16 integers, considering their ascending order.   Problems 3 (Before Optimization): A fair coin is tossed 50 times, what is the probability of getting heads at least 25 times? Problems 3 (After Optimization): A fair coin is tossed 50 times; what is the probability of obtaining heads at least 25 times, and can you also calculate the expected number, variance, and standard deviation of heads while determining the likelihood that the total number of heads exceeds 30?
@@
ScaleQuest Example 1 QUESTION: My orthogonal projection matrix $P$ projects onto the plane with equation $x+y+z=0.$ Find the matrix $P$.   SOLUTION: To find the orthogonal projection matrix $P$ that projects onto the plane with the equation $x+y+z=0$, we can follow these steps: 1. Identify the normal vector to the plane:
@@ 2. Express the projection matrix in terms of the normal vector:
ScaleQuest Example 3 QUESTION: Compute $\sum_{n=1}^{\infty}\frac{4^{n}+5^{n}+6^{n}}{12^{n}}.$   SOLUTION: To compute the sum $\sum_{n=1}^{\infty}\frac{4^{n}+5^{n}+6^{n}}{12^{n}}$, we can split the sum into three separate sums: $\sum_{n=1}^{\infty}\frac{4^{n}+5^{n}+6^{n}}{12^{n}}=\sum_{n=1}^{\infty}\frac{4^{n}}{12^{n}}+\sum_{n=1}^{\infty}\frac{5^{n}}{12^{n}}+\sum_{n=1}^{\infty}\frac{6^{n}}{12^{n}}.$ Each of these sums is a geometric series. Let’s evaluate each one separately. 1. For the first sum $\sum_{n=1}^{\infty}\frac{4^{n}}{12^{n}}$: $\sum_{n=1}^{\infty}\frac{4^{n}}{12^{n}}=\sum_{n=1}^{\infty}\left(\frac{4}{12}\right)^{n}=\sum_{n=1}^{\infty}\left(\frac{1}{3}\right)^{n}.$ The sum of an infinite geometric series $\sum_{n=1}^{\infty}ar^{n-1}$ is given by $\frac{a}{1-r}$ for $|r|<1$. Here, $a=\frac{1}{3}$ and $r=\frac{1}{3}$, so: $\sum_{n=1}^{\infty}\left(\frac{1}{3}\right)^{n}=\frac{\frac{1}{3}}{1-\frac{1}{3}}=\frac{\frac{1}{3}}{\frac{2}{3}}=\frac{1}{2}.$ 2. For the second sum $\sum_{n=1}^{\infty}\frac{5^{n}}{12^{n}}$: $\sum_{n=1}^{\infty}\frac{5^{n}}{12^{n}}=\sum_{n=1}^{\infty}\left(\frac{5}{12}\right)^{n}.$ Here, $a=\frac{5}{12}$ and $r=\frac{5}{12}$, so: $\sum_{n=1}^{\infty}\left(\frac{5}{12}\right)^{n}=\frac{\frac{5}{12}}{1-\frac{5}{12}}=\frac{\frac{5}{12}}{\frac{7}{12}}=\frac{5}{7}.$ 3. For the third sum $\sum_{n=1}^{\infty}\frac{6^{n}}{12^{n}}$: $\sum_{n=1}^{\infty}\frac{6^{n}}{12^{n}}=\sum_{n=1}^{\infty}\left(\frac{6}{12}\right)^{n}=\sum_{n=1}^{\infty}\left(\frac{1}{2}\right)^{n}.$ Here, $a=\frac{1}{2}$ and $r=\frac{1}{2}$, so: $\sum_{n=1}^{\infty}\left(\frac{1}{2}\right)^{n}=\frac{\frac{1}{2}}{1-\frac{1}{2}}=\frac{\frac{1}{2}}{\frac{1}{2}}=1.$ Adding these three results together, we get: $\sum_{n=1}^{\infty}\frac{4^{n}+5^{n}+6^{n}}{12^{n}}=\frac{1}{2}+\frac{5}{7}+1.$ To add these fractions, we need a common denominator. The least common multiple of 2, 7, and 1 is 14. So we convert each fraction: $\frac{1}{2}=\frac{7}{14},\quad\frac{5}{7}=\frac{10}{14},\quad 1=\frac{14}{14}.$ Adding these fractions together, we get: $\frac{7}{14}+\frac{10}{14}+\frac{14}{14}=\frac{7+10+14}{14}=\frac{31}{14}.$ Thus, the sum is: $\boxed{\frac{31}{14}}.$

### v2

@@ Appendix B Additional Data Statistics
Appendix C ScaleQuest for Broader tasks
C.1 Long Chain-of-Thought Mathematical Reasoning
DeepSeek-R1 (Guo et al., 2025) demonstrated the effectiveness of Long Chain-of-Thought (CoT) reasoning in tackling challenging mathematical reasoning tasks. In this part, we further investigate the capability of ScaleQuest to generate high-quality long CoT data.
Settings
We utilize the latest Qwen2.5-Math-7B model (Yang et al., 2024b) for question fine-tuning and preference optimization. To construct long CoT data, we select DeepSeek-R1-Distill-Qwen-7B as the response generator. The rest of the pipeline remains unchanged from prior work. In total, we synthesized 150K Long Chain-of-Thought training samples.
Results
We then fine-tuned DeepSeek-R1-Distill-Qwen-7B on this dataset, which also serves as the teacher model during data synthesis. Table 8 presents the results, showing that the fine-tuned model outperforms its teacher, demonstrating a strong potential for self-improvement.
Model GSM8k MATH500 AIME24 AMC25 Average
DS-R1-Distill-Qwen-7B 91.5 90.2 43.3 75.0 75.0
DS-R1-Distill-Qwen-7B-ScaleQuest 93.0 91.0 53.3 90.0 81.8
Table 8: Results of ScaleQuest-LongCoT.
C.2 Code Reasoning
We also extend our ScaleQuest method to the Code Reasoning Task as a simple validation. We made the following modifications to adapt to the code reasoning task:
Settings
We choose DeepSeek-Coder-7B-Instruct (Guo et al., 2024) and Qwen2.5-Coder-7B-Instruct (Hui et al., 2024) as two problem-solving models to perform question fine-tuning on 20K questions randomly sampled from CodeFeedBack (Zheng et al., 2024). For Question Preference Optimization, we also focus on solvability and difficulty, making slight modifications to the prompts based on the code reasoning task. Our evaluation cover HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and BigCodeBench (Zhuo et al., 2024), using the same evaluation script as Qwen2.5-Coder. We report pass@1 results using greedy search.
Model # Samples (K) HumanEval MBPP BigCodeBench Average
Qwen2.5-Coder-CFB 156 79.3 77.2 35.6 64.0
Qwen2.5-Coder-CFB-Aug 156 84.1 84.7 39.0 69.3
Qwen2.5-Coder-ScaleQuest 156 86.6 83.1 40.0 69.9
Table 9: Results of ScaleQuest in Code Reasoning Task. All results are based on Qwen2.5-Coder-7B-Base. CFB refers to the CodeFeedBack-Filtered Dataset. We augment the responses for the problems in CodeFeedback-Filtered using Qwen2.5-Coder-7B-Instruct with reward filtering, creating a new dataset referred to as CFB-Aug.
Results
The results are presented in Table 9. Compared to the widely used refined version of CodeFeedback, namely CodeFeedback-Filtered, our generated data outperforms it, with an average improvement of 5.9 across the three baselines. Additionally, we enhance the Response portion of CodeFeedback-Filtered using Qwen2.5-Coder-7B-Instruct, and the results indicate that our generated questions are of higher quality. This further demonstrates the effectiveness of the ScaleQuest method.
Appendix D Insights behind model selection
In our works, we use multiple models, e.g., DSMath-7B-RL, Qwen2-Math-7B-Ins, GPT-4o-mini, and DSMath-7B-Base, which may cause confusion for model selection. In response, we also supplement our approach with a simpler setup. We use Qwen2-Math-7B-Ins for training question generators, constructing optimization data for QPO, and performing solvability & difficulty filtering, as well as for response generation. For reward filtering, InternLM-7B-Reward remained unchanged. The results, as shown in Figure 11 (ScaleQuest-Simple result), indicate that our approach continues to demonstrate superior performance compared to existing datasets. Additionally, we summarize these insights on model selection for domain adaptation:
  • Selection of base model for training question generator: The self-synthesis generation paradigm heavily relies on the inherent knowledge of the problem-solving model itself (Xu et al., 2024). Therefore, a domain-specific model is essential. For example, Qwen2-Math-Ins is suitable for mathematical reasoning, while Qwen2.5-Coder-Ins fits well for code reasoning. Furthermore, using multiple question generators often leads to more diverse and higher-quality questions (as discussed in section 3.3).
  • Selection of model for constructing optimization data: Well-aligned, general-purpose models, such as Llama3.1-70B and GPT-4o-mini, tend to perform better than domain-specific models, as illustrated in Figure 4.
  • Selection of Response Generation Model & Reward Model: These can be selected based on their performance on the corresponding mathematical tasks.
We believe that the methodology and the experience in selecting models are always more critical than the chosen models themselves. With the continuous advancements in the open-source community, we are confident that stronger models will undoubtedly produce even better datasets when applying our approach.
Appendix E Additional Experiments
E.1 Evaluation Results on More Out-of-Domain (OOD) Benchmarks
In addition to College Math and Olympiad Bench, we include two additional OOD benchmarks: GSM-Hard (Gao et al., 2023) and MathChat (Liang et al., 2024). GSM-Hard is constructed by modifying the questions in GSM8K, replacing the numbers with larger, less common ones. From MathChat, we select two problem-solving tasks: follow-up QA and error correction. The results are summarized in Table 10. In more fine-grained OOD evaluations, our model continues to perform on par with Qwen2-Math-7B-Ins, further demonstrating our ScaleQuest Model’s generalization capability and highlighting the generated data’s robustness.
Model GSM-Hard Follow-up QA Error Correction Average
R1 R2 R3
Qwen2-Math-7B-Instruct 68.3 89.5 62.4 53.5 89.9 72.7
Qwen2-Math-7B-ScaleQuest 66.3 89.7 61.7 53.5 91.1 72.5
Table 10: The comparison between Qwen2-Math-7B-Ins and the ScaleQuest Model on GSM-Hard and MathChat. We choose Follow-up QA and Error Correction from MathChat for evaluation in problem-solving. R1, R2, and R3 represent different rounds in Follow-up QA.
E.2 Comparison Under Equal Training Data Volume
In the right panel of Figure 1, we plot the scaling trends of model performance with increasing data volume, showcasing the superiority of the ScaleQuest method when using the same amount of data. To further ensure a fair comparison, we randomly sample the same number of training examples from open-source datasets for training. Specifically, we sample 400K examples from MetaMath, DART-Math, NuminaMath, and our dataset (for MetaMath, which contains 395K examples in total, all samples are used). The results are presented in Table 11. We observe that with the same amount of training data, our dataset demonstrates significantly higher instruction tuning effectiveness compared to other datasets.
Model # Samples (K) GSM8K MATH College Math Olympiad Bench Average
Qwen2-Math-7B-MetaMath 395 84.3 48.6 40.5 15.6 47.3
Qwen2-Math-7B-DART-Math 400 88.6 58.2 45.2 22.8 53.7
Qwen2-Math-7B-NuminaMath 400 82.0 65.8 44.9 29.2 55.5
Qwen2-Math-7B-ScaleQuest 400 90.6 71.6 50.2 36.2 62.1
Qwen2-Math-7B-ScaleQuest-Simple 400 89.4 69.9 48.8 33.6 60.4
Table 11: Results on four mathematical reasoning benchmarks. All results are based on Qwen2-Math-7B-Base. ScaleQuest-Simple is a simplified version that only utilizes Qwen2-Math-7B-Ins for QFT, QPO, and question filtering, and InternLM-7B-Reward for reward filtering.
Model GSM8K MATH College Math Olympiad Bench Average
Mistral-7B-MetaMath-Aug 77.0 34.1 18.6 8.6 34.6
Mistral-7B-OrcaMath-Aug 84.4 31.6 20.9 8.2 36.3
Mistral-7B-NumiMath-Aug 79.5 62.8 40.4 30.4 53.3
Mistral-7B-ScaleQuest 88.5 62.9 43.5 28.8 55.9
Llama3-8B-MetaMath-Aug 77.6 33.1 20.6 9.2 35.1
Llama3-8B-OrcaMath-Aug 83.2 32.6 19.4 8.6 36.0
Llama3-8B-NumiMath-Aug 79.1 62.9 39.3 25.4 51.7
Llama3-8B-ScaleQuest 87.9 64.4 42.8 25.3 55.1
Qwen2-Math-7B-MetaMath-Aug 88.5 68.5 47.1 33.0 59.3
Qwen2-Math-7B-OrcaMath-Aug 89.3 68.3 46.6 31.9 59.0
Qwen2-Math-7B-NumiMath-Aug 89.5 72.6 49.5 36.3 62.0
Qwen2-Math-7B-ScaleQuest 89.7 73.4 50.0 38.5 62.9
Table 12: Additional results of Table 2 on the other base models. All responses are generated using Qwen2-Math-7B-Instruct with the same reward filtering process. For baseline datasets, “-Aug” indicates that the responses have been enhanced.
E.3 Training Data Volume on QPO
QPO is designed to enhance the solvability and difficulty of the question generator. We investigate the impact of training data volume by using GPT-4o-mini as the optimization model. The training data volume is controlled at 5K, 10K, 15K, 20K, and 40K, with Qwen2-Math-7B-QFT serving as the base model. We evaluate the performance of the trained question generator in terms of solvability and difficulty. The results are shown in Figure 9. As the amount of training data increases, both the solvable rate and difficulty of the questions generated by the question generator improve, gradually converging around 20K training examples. We believe that maintaining the training data at approximately 10K represents a more suitable balance between training cost and model performance.
Figure 9: Performance of QPO in different training data volume. The evaluation covers the solvable ratio and difficulty score, following the same evaluation procedure as in Figure 5.
E.4 Additional Results Based on Different Base Models
We have supplemented Table 2 with the results for the other three base models, as shown in Table 12. Under the same response generation process, our approach consistently outperforms existing datasets across all four base models, further demonstrating the superiority of our method.
Appendix F Prompt Template
Prompts for Problem Solvability Optimization Please act as a professional math teacher.
Your goal is to create high quality math word problems to help students learn math.
You will be given a math question. Please optimize the Given Question and follow the instructions.
To achieve the goal, please follow the steps:
# Please check that the given question is a math question and write detailed solution to the Given Question.
# Based on the problem-solving process, double check the question is solvable.
# If you feel that the given question is not a meaningful math question, rewrite one that makes sense to you. Otherwise, modify the Given question according to your checking comment to ensure it is solvable and of high quality.
# If the question can be solved with just a few simple thinking processes, you can rewrite it to explicitly request multiple-step reasoning.
You have five principles to do this:
# Ensure the optimized question only asks for one thing, be reasonable and solvable, be based on the Given Question (if possible), and can be answered with only a number (float or integer). For example, DO NOT ask, ‘what is the amount of A, B and C?’.
# Ensure the optimized question is in line with common sense of life. For example, the amount someone has or pays must be a positive number, and the number of people must be an integer.
# Ensure your student can answer the optimized question without the given question. If you want to use some numbers, conditions or background in the given question, please restate them to ensure no information is omitted in your optimized question.
# Please DO NOT include solution in your question.
Given Question: problem
Your output should be in the following format:
CREATED QUESTION: [your created question]
VERIFICATION AND MODIFICATION: [solve the question step-by-step and modify it to follow all principles]
FINAL QUESTION: [your final created question]
Figure 10: Prompts used to optimize the solvability of questions for QPO Training.
Prompts for Problem Difficulty Optimization You are an Math Problem Rewriter that rewrites the given #Problem# into a more complex version.
Please follow the steps below to rewrite the given "#Problem#" into a more complex version.
@@ Appendix F Prompt Template
Step 1: Please read the "#Problem#" carefully and list all the possible methods to make this problem more complex (to make it a bit harder for well-known AI assistants such as ChatGPT and GPT4 to handle). Note that the problem itself might be erroneous, and you need to first correct the errors within it.
Step 2: Please create a comprehensive plan based on the #Methods List# generated in Step 1 to make the #Problem# more complex. The plan should include several methods from the #Methods List#.
Step 3: Please execute the plan step by step and provide the #Rewritten Problem#. #Rewritten Problem# can only add 10 to 20 words into the "#Problem#".
Step 4: Please carefully review the #Rewritten Problem# and identify any unreasonable parts. Ensure that the #Rewritten Problem# is only a more complex version of the #Problem#. Just provide the #Finally Rewritten Problem# without any explanation and step-by-step reasoning guidance.
Please reply strictly in the following format:
Step 1 #Methods List#:
Step 2 #Plan#:
Step 3 #Rewritten Problem#:
Step 4 #Finally Rewritten Problem#:
#Problem#: Problem
Figure 11: Prompts used to optimize the difficulty of questions for QPO Training.
Prompts for Problem Solvability Check Please act as a professional math teacher.
Your goal is to determine if the given problem is a valuable math problem. You need to consider two aspects:
1. The given problem is a math problem.
@@
2. The given math problem can be solved based on the conditions provided in the problem (You can first try to solve it and then judge its solvability).
Please reason step by step and conclude with either ‘Yes’ or ‘No’.
Given Problem: Problem
Figure 12: Prompts used to check the solvability of questions.
Prompts for Difficulty Classification # Instruction
You first need to identify the given user intent and then label the difficulty level of the user query based on the content of the user query.
## User Query
Input
## Output Format
Given the user query, in your output, you first need to identify the user intent and the knowledge needed to solve the task in the user query.
Then, rate the difficulty level of the user query as very easy, easy, medium, hard, or very hard.
Now, please output the user intent and difficulty level below in a json format by filling in the placeholders in []:
‘‘intent’’: ‘‘The user wants to [....]’’,
‘‘knowledge’’: ‘‘To solve this problem, the models need to know [....]’’,
‘‘difficulty’’: ‘‘[very easy/easy/medium/hard/very hard]’’
@@
Prompts for Topic Classification As a mathematics education specialist, please analyze the topics of the provided question and its answer.
Specific requirements are as follows:
1. You should identify and categorize the main mathematical topics involved in the problem. If knowledge from non-mathematical fields is used, it is classified into Others - xxx, such as Others - Problem Context.
2. You should put your final answer between $<$TOPIC$>$ and $<$/TOPIC$>$.
Question: Compute $\cos 330^{\circ}$.
Answer: We know that $330^{\circ}=360^{\circ}-30^{\circ}$.
Since $\cos(360^{\circ}-\theta)=\cos\theta$ for all angles $\theta$,
we have $\cos 330^{\circ}=\cos 30^{\circ}$.
Since $\cos 30^{\circ}=\frac{\sqrt{3}}{2}$,
we can conclude that $\cos 330^{\circ}=\boxed{\frac{\sqrt{3}}{2}}$.
Analysis: $<$TOPIC$>$Trigonometry - Cosine Function$<$/TOPIC$>$
Question: Question
Answer: Answer
Analysis:
@@
Examples for Solvability Optimization Problems 1 (Before Optimization):
There are 10 survivors in an emergency room. Each survivor is either a child, a woman, or a man. If there are 4 men and 3 times as many women as men, how many children are there?
Problems 1 (After Optimization):
There are 10 survivors in an emergency room. Each survivor is either a child, a woman, or a man. If there are 4 men and an equal number of women as men, how many children are there?
Problems 2 (Before Optimization):
How many sides does a polygon have if it is a regular polygon?
Problems 2 (After Optimization):
How many sides does a regular polygon have if each interior angle is 120 degrees?
Problems 3 (Before Optimization):
Find the sum of the first three terms of this series.
Problems 3 (After Optimization):
Calculate the sum of the first three terms of the arithmetic series where the first term is 5 and the common difference is 3.
@@
Examples for Difficulty Optimization Problems 1 (Before Optimization):
How many 4-digit positive integers are there?
Problems 1 (After Optimization):
How many 4-digit positive integers can be formed using non-repeating digits where the sum of these digits must be even, and the integers fall within the range of 1000 to $9$?
Problems 2 (Before Optimization):
The average of 15 numbers is 32. An additional number is then added to the list, and the new average of the 16 numbers is 34. What number was added to the list?
Problems 2 (After Optimization):
The average of 15 positive integers is 32, but one integer fluctuates to 30 before adding a new number. After adding this new number, the average of the 16 integers becomes 34. Calculate the added number and find the standard deviation of all 16 integers, considering their ascending order.
Problems 3 (Before Optimization):
A fair coin is tossed 50 times, what is the probability of getting heads at least 25 times?
Problems 3 (After Optimization):
A fair coin is tossed 50 times; what is the probability of obtaining heads at least 25 times, and can you also calculate the expected number, variance, and standard deviation of heads while determining the likelihood that the total number of heads exceeds 30?
@@
ScaleQuest Data Example 1 QUESTION:
My orthogonal projection matrix $P$ projects onto the plane with equation
$x+y+z=0.$ Find the matrix $P$.
  SOLUTION:
To find the orthogonal projection matrix $P$ that projects onto the plane with the equation $x+y+z=0$, we can follow these steps:
1. Identify the normal vector to the plane:
@@ 2. Express the projection matrix in terms of the normal vector:
ScaleQuest Data Example 2 QUESTION:
Suppose $a,b$ are positive real numbers such that $a+b=1$.
(a) minimize $\frac{1}{a}+\frac{1}{b}$; (b) minimize $a^{2}+b^{2}$.
  SOLUTION:
To solve these problems, we will use the method of Lagrange multipliers and the AM-HM inequality.
@@
ScaleQuest Data Example 3 QUESTION:
Compute
$\sum_{n=1}^{\infty}\frac{4^{n}+5^{n}+6^{n}}{12^{n}}.$   SOLUTION:
To compute the sum $\sum_{n=1}^{\infty}\frac{4^{n}+5^{n}+6^{n}}{12^{n}}$, we can split the sum into three separate sums:
$\sum_{n=1}^{\infty}\frac{4^{n}+5^{n}+6^{n}}{12^{n}}=\sum_{n=1}^{\infty}\frac{4^{n}}{12^{n}}+\sum_{n=1}^{\infty}\frac{5^{n}}{12^{n}}+\sum_{n=1}^{\infty}\frac{6^{n}}{12^{n}}.$ Each of these sums is a geometric series. Let’s evaluate each one separately. 1. For the first sum $\sum_{n=1}^{\infty}\frac{4^{n}}{12^{n}}$: $\sum_{n=1}^{\infty}\frac{4^{n}}{12^{n}}=\sum_{n=1}^{\infty}\left(\frac{4}{12}\right)^{n}=\sum_{n=1}^{\infty}\left(\frac{1}{3}\right)^{n}.$ The sum of an infinite geometric series $\sum_{n=1}^{\infty}ar^{n-1}$ is given by $\frac{a}{1-r}$ for $|r|<1$. Here, $a=\frac{1}{3}$ and $r=\frac{1}{3}$, so: $\sum_{n=1}^{\infty}\left(\frac{1}{3}\right)^{n}=\frac{\frac{1}{3}}{1-\frac{1}{3}}=\frac{\frac{1}{3}}{\frac{2}{3}}=\frac{1}{2}.$ 2. For the second sum $\sum_{n=1}^{\infty}\frac{5^{n}}{12^{n}}$: $\sum_{n=1}^{\infty}\frac{5^{n}}{12^{n}}=\sum_{n=1}^{\infty}\left(\frac{5}{12}\right)^{n}.$ Here, $a=\frac{5}{12}$ and $r=\frac{5}{12}$, so: $\sum_{n=1}^{\infty}\left(\frac{5}{12}\right)^{n}=\frac{\frac{5}{12}}{1-\frac{5}{12}}=\frac{\frac{5}{12}}{\frac{7}{12}}=\frac{5}{7}.$ 3. For the third sum $\sum_{n=1}^{\infty}\frac{6^{n}}{12^{n}}$: $\sum_{n=1}^{\infty}\frac{6^{n}}{12^{n}}=\sum_{n=1}^{\infty}\left(\frac{6}{12}\right)^{n}=\sum_{n=1}^{\infty}\left(\frac{1}{2}\right)^{n}.$ Here, $a=\frac{1}{2}$ and $r=\frac{1}{2}$, so: $\sum_{n=1}^{\infty}\left(\frac{1}{2}\right)^{n}=\frac{\frac{1}{2}}{1-\frac{1}{2}}=\frac{\frac{1}{2}}{\frac{1}{2}}=1.$ Adding these three results together, we get: $\sum_{n=1}^{\infty}\frac{4^{n}+5^{n}+6^{n}}{12^{n}}=\frac{1}{2}+\frac{5}{7}+1.$ To add these fractions, we need a common denominator. The least common multiple of 2, 7, and 1 is 14. So we convert each fraction: $\frac{1}{2}=\frac{7}{14},\quad\frac{5}{7}=\frac{10}{14},\quad 1=\frac{14}{14}.$ Adding these fractions together, we get: $\frac{7}{14}+\frac{10}{14}+\frac{14}{14}=\frac{7+10+14}{14}=\frac{31}{14}.$ Thus, the sum is: $\boxed{\frac{31}{14}}.$